At the beginning of 2014, Robotics Inc. acquired a manufacturing facility for $14.0 million. $11.0 million of the purchase price was allocated to the building. Depreciation for 2014 and 2015 was calculated using the straight-line method, a 20-year useful life, and a $3.0 million residual value. In 2016 the company switched to the double-declining-balance depreciation method What is depreciation on the building for 2016? (Do not round intermediate calculations. Enter your answer in whole dollars.) Depreciation .130,000
Answer #1

Depreciation as per straight line= (original cost-salvage value)/number of years

Depreciation as per straight line={11000000-3000000)/20

Depreciation as per straight line= 400000

Depreciation rate= 400000/8000000*100= 5%

Depreciation rate as per double declining method= 2*5= 10%

Depreciation as per double declining method= (11000000-400000-400000)*0.10

Depreciation as per double declining method= 10200000*0.10

Depreciation as per double declining method= 1020000